OpenStudy (superhelp101):

How many milliliters of sodium metal, with a density of 0.97 g/mL, would be needed to produce 53.2 grams of hydrogen gas in the single-replacement reaction below? Show all steps of your calculation as well as the final answer. Na + H2O → NaOH + H2

OpenStudy (superhelp101):

can u plz explain to me how to solve this i am really confused?

OpenStudy (abmon98):

1) Balance Write down the number of atoms that you have on each side of the equation. Na + H2O → NaOH + H2 On the reactant Side: Na:1 H:2 O:1 On the Product Side : Na:1 H:3 O:1 Add a coefficient of 2 to H2O Na + 2H2O →NaOH + H2 On the reactant Side: Na:1 H:4 O:2 On the Product Side: Na:1 H:3 O:1 Add a coeffcient of 2 to both Na and NaOH 2Na + 2H2O →2NaOH + H2 On the reactant Side: Na:2 H:4 O:2 On the Product Side: Na:2 H:4 O:2 Number of Moles=Mass(g)/Molar Mass(g/mol) 53.2/2=26.6 Look at the Mole Ratio of Na:H2 2:1 x:26.6 26.6*2/1=53.2 moles Mass(g)=Molar Mass(g/mol)*Number of moles 53.2*23.0=1223.6 grams of Sodium Density=Mass/Volume 1223.6/V=0.97 0.97V=1223.6 V=1261.4 ml

OpenStudy (superhelp101):

WOW thank for explaining this to me!!